# Do spectral cues matter in contrast-based graph self-supervised learning?

### v1

@@ Abstract
The recent surge in contrast-based graph self-supervised learning has prominently featured an intensified exploration of spectral cues. However, an intriguing paradox emerges, as methods grounded in seemingly conflicting assumptions or heuristic approaches regarding the spectral domain demonstrate notable enhancements in learning performance. This paradox prompts a critical inquiry into the genuine contribution of spectral information to contrast-based graph self-supervised learning. This study undertakes an extensive investigation into this inquiry, conducting a thorough study of the relationship between spectral characteristics and the learning outcomes of contemporary methodologies. Based on this analysis, we claim that the effectiveness and significance of spectral information need to be questioned. Instead, we revisit simple edge perturbation: random edge dropping designed for node-level self-supervised learning and random edge adding intended for graph-level self-supervised learning. Compelling evidence is presented that these simple yet effective strategies consistently yield superior performance while demanding significantly fewer computational resources compared to all prior spectral augmentation methods. The proposed insights represent a significant leap forward in the field, potentially reshaping the understanding and implementation of graph self-supervised learning.
@@ 1 Introduction
In recent years, graph learning has emerged as a powerhouse for handling complex data relationships in multiple fields, offering vast potential and value, particularly in domains such as data mining [10], computer vision [35], network analysis [4], and bioinformatics [12]. However, limited labels make graph learning challenging to apply in real-world scenarios. Inspired by the great success of Self-Supervised Learning (SSL) in other domains [6, 3], Graph Self-Supervised Learning (Graph SSL) has made rapid progress and has shown promise by achieving state-of-the-art performance on many tasks [34], where Contrast-based Graph SSL (CG-SSL) are most dominate [20]. This type of method is grounded in the concept of mutual information (MI) maximization. The primary goal is to maximize the estimated MI between augmented instances of the same object, such as nodes, subgraphs, or entire graphs. Among the new developments in CG-SSL, approaches inspired by graph spectral methods have garnered significant attention. A prevalent conviction is that spectral information, including the eigenvalues and eigenvectors of the graph’s Laplacian, plays a crucial role in enhancing the efficacy of CG-SSL [18, 14, 16, 36].
@@
Several methodologies heavily emphasize the utilization of spectral information while implementing sampling or sparsification procedures to enhance learning capabilities. However, there seems no consensus on the effectiveness of spectral information can be made. SpCo [18] introduces the general graph augmentation (GAME) rule, which suggests that the difference in high-frequency parts between augmented graphs should be larger than that of low-frequency parts. SPAN [16] contends that effective topology augmentation in Graph Contrastive Learning (GCL) should prioritize perturbing sensitive edges that have a substantial impact on the graph spectrum. Therefore, a principled augmentation method is designed by directly maximizing spectral change with a certain perturbation budget. GASSER [36] selectively perturbs graph structures based on spectral cues to better maintain the required invariance for contrastive learning frameworks. A contradiction emerges among these related works on spectral augmentation: SPAN advocates for maximizing the disparity between augmented graphs, while SpCo and GASSER argue for the preservation of specific spectral components during augmentation. Moreover, SpCo underscores the task-agnostic significance of retaining information from low-frequency components, whereas GASSER suggests that augmentations should selectively preserve task-relevant frequency components while cautiously perturbing task-irrelevant ones. The consistent performance gain derived from opposing methodical designs naturally raises our concern:
   •  Does spectral cues really matter in contrast-based graph SSL?
Given the question, this study aims to critically evaluate the effectiveness and significance of spectral augmentation in contrast-based graph SSL frameworks (CG-SSL). With evidence-supported claims and findings in the following sections, we can give a negative answer to the question above: No, they are not very effective and we don’t really need them. To be specific, we find that spectral augmentation does not significantly contribute to the learning efficacy while more straightforward edge perturbations are already good enough for contrast-based graph SSL. We manage to elaborate on our conclusion through a series of studies carried out in the following efforts:
@@
    In Sec. 4, we explore the dependency of spectral augmentation effectiveness on the depth of the network, positing that shallower networks with fewer convolutional layers perform better but demonstrate diminished benefits from spectral changes.
@@
    In Sec 5 We claim that simple edge perturbation techniques, like adding edges to or dropping edges from the graph, not only compete well but often outperform spectral augmentations, without any significant help from spectral cues. To support this,
    (a) In Sec. 6, overall model performance on test accuracy with four state-of-the-art frameworks on both node- and graph-level classification tasks support the superiority of simple edge perturbation. (b) Studies in Sec. 7.1 reveal the indistinguishability between the average spectrum of augmented graphs from edge perturbation with optimal parameters on different datasets, no matter how different that of original graphs is, indicating GNN encoders can hardly learn spectral information from augmented graphs. That is to say, edge perturbations can not benefit from spectral cues. (c) In Sec. 7.2, we analyze the effectiveness of state-of-the-art spectral augmentation baseline (i.e., SPAN) by perturbing edges to alter the spectral characteristics of augmented graphs from simple edge perturbation augmentation and examining the impact on model performance. As it turns out, the results show no performance degradation, indicating the spectral information contained in the augmentation is not significant to the model performance. (d) In Appendix E.2, statistical analysis is carried out to argue that the major reason edge perturbation works well is not because of the spectral information as they are not the key factor on model performance.
@@ 2 Related work
Graph Contrastive Learning. Contrast-based Graph Self-Supervised (CG-SSL) learning alleviates the limitations of supervised learning, which heavily depends on labeled data and often suffers from limited generalization [19]. This makes it a promising approach for real-world applications where labeled data is scarce. In general, CG-SSL can be categorized into Graph Contrastive Learning (GCL) and Graph Predictive Learning [34], here we focus our discussion on the former. GCL applies a variety of augmentations to the training graph to obtain augmented views. These augmented views, which are derived from the same original graph, are treated as positive sample pairs or sets. The key objective of GCL is to maximize the mutual information between these views to learn robust and invariant representations. However, directly computing the mutual information of graph representations is challenging. Hence, in practice, GCL frameworks aim to maximize the lower bound of mutual information using different estimators such as InfoNCE [9], Jensen-Shannon [22], and Donsker-Varadhan [1]. For instance, frameworks like GRACE [42], GCC [25], and GCA [43] utilize the InfoNCE estimator as their objective function. On the other hand, MVGRL [11] and InfoGraph [28] adopt the Jensen-Shannon estimator.
Graph Augmentations in GCL. Beyond the choice of object functions, another crucial aspect of GCL is the selection of augmentation techniques. Early work by [42] and [37] introduced several heuristic domain-agnostic graph augmentation for CG-SSL, such as edge perturbation, attribute masking, and subgraph sampling. These straightforward and effective methods have been widely adopted in subsequent CG-SSL frameworks due to their demonstrated success [29, 38]. However, these domain-agnostic graph augmentations often lack interpretability, making it difficult to understand the exact impact of these augmentations on the graph structure and learning outcomes. To address this issue, MVGRL [11] introduces graph diffusion as an augmentation strategy, where the original graph provides local structural information and the diffused graph offers global context. MVGRL demonstrates experimentally that by optimizing for consistency between node representations from these two perspectives, it’s possible to obtain representations that encode both local and global structural information. Moreover, SpCo [18], GASSER [36], and SPAN [16] provide design principles for graph augmentation based on spectral graph theory, further elaborating on how graph augmentation can help CG-SSL. However, our explorations show that these methods are unable to consistently outperform heuristic graph augmentations such as edge perturbation (DropEdge or AddEdge) in terms of performance under fair comparisons, and thus the design principles of graph augmentation still require further validation.
@@ 3 Preliminary study
Contrast-based graph self-supervised learning framework.
CG-SSL captures invariant features of a graph by generating multiple views (typically two) through augmentations and then maximizing the mutual information between these views [34]. This approach is ultimately used to improve performance on various downstream tasks. Following previous work [33, 19, 34], we first denote the generic form of the augmentation $\mathcal{T}$ and objective functions $\mathcal{L}_{cl}$ of graph contrastive learning. Given a graph $\mathcal{G}=(\mathbf{A},\mathbf{X})$ with adjacency matrix $\mathbf{A}$ and feature matrix $\mathbf{X}$, the augmentation is defined as the transformation function $\mathcal{T}$. In this paper, we are mainly concerned with topological augmentation, in which feature matrix $\mathbf{X}$ remains intact:
@@
In this paper, we utilize four prominent CG-SSL frameworks to study the effect of spectral: MVGRL, GRACE, BGRL, and G-BT. MVGRL introduces graph diffusion as augmentation, while the other three frameworks use edge perturbation as augmentation. Each framework employs different strategies for its contrastive loss functions. MVGRL and GRACE use the Jensen-Shannon and InfoNCE estimators as object functions, respectively. In contrast, BGRL and G-BT adopt the BYOL loss [7] and Barlow Twins loss [39], which are designed to maximize the agreement between the augmented views without relying on negative samples. A more detailed description of the loss function can be found in the Appendix D.
Graph spectrum and application of spectral augmentation.
We follow the standard definition of graph spectrum in this study, details of which can be found in Appendix C. Among various augmentation strategies proposed to enhance the robustness and generalization of graph neural networks, spectral augmentation has been considered a promising avenue [16, 18, 2, 36]. Spectral augmentation typically involves implicit modifications to the eigenvalues of the graph Laplacian, aimed at enhancing model performance by encouraging invariance to certain spectral properties. Among them, SPAN achieved state-of-the-art performance in both node classification and graph classification. In short, SPAN elaborates two augmentation functions, $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, where $\mathcal{T}_{1}$ maximizes the spectral norm in one view, and $\mathcal{T}_{2}$ minimizes it in the other view. Subsequently, these two augmentations are implemented in the four CG-SSL frameworks mentioned above (Strict definition in Appendix C). The paradigm used by SPAN aims to allow the GNN encoder to focus on robust spectral components and ignore the sensitive edges that can change the spectral drastically when perturbed.
@@ 4 Limitations of spectral augmentations
Limitations of shallow GNN encoders in capturing spectral information.
Multiple previous studies indicate that shallow, rather than deep, GNN encoders can be effective in graph self-supervised learning. This might be the result of overfitting commonly witnessed in standard GNN tasks. We have also carried out many empirical studies with a range of CG-SSL frameworks and augmentations to support this idea in contrast-based graph SSL. As the most commonly applied GNN encoder in CG-SSL [37, 38, 8, 17], an empirical study on the relationship between the depth of GCN encoder and learning performance is conducted and results are presented in Fig. 1. From that, we can conclude that shallow GCN encoders with 1 or 2 layers usually have the best performance. Note that this tendency is not very clear on graph-level tasks because the embedding of the graph from all layers will be concatenated together to perform prediction. It indicates a deep encoder has theoretically better expressive power than any encoder shallower than it. Therefore, still better performance of GCN encoders with 1 or 2 layers implies that any more layers are unnecessary and might hurt the quality of the learned representation of the graph.
@@
Figure 1: Accuracy of CG-SSL v.s. number of GCN layers on node and graph classification on representative datasets. (a) G-BT on node classification. (b) MVGRL on node classification. (c) G-BT on graph classification. (d) MVGRL on graph classification. We choose two representative datasets for each task, i.e. Cora and CiteSeer for the node-level and PROTEINS and IMDB-BINARY for the graph-level classification. Definition of the accuracy of CG-SSL, details of datasets and other experimental settings are mentioned in Section 6.1.
@@
By design, most GNN encoders primarily aggregate local neighborhood information through their layered structure, where each layer extends the receptive field by one hop. The depth of a GNN critically determines its ability to integrate information from various parts of the graph. With only a limited number of layers, a GNN’s receptive field is restricted to immediate neighborhoods (e.g., 1-hop or 2-hop distances). This limitation severely constrains the network’s ability to assimilate and leverage broader graph topologies or global features that are essential for encoding the spectral properties of the graph.
Limited implications for spectral augmentation in CG-SSL.
@@
Given the limitations of shallow GNNs in capturing spectral information, the utility of spectral augmentation techniques in graph self-supervised learning settings warrants scrutiny. Spectral augmentation involves modifying the spectral components (e.g., eigenvalues and eigenvectors) of a graph to enrich the training process or to create diverse samples for enhancing learning robustness. However, if the primary encoder’s architecture—specifically, a shallow GNN—is intrinsically limited in its ability to perceive and process spectral properties, then the benefits of such augmentations are likely to be minimal.

### v2

# Rethinking Spectral Augmentation for Contrast-based Graph Self-Supervised Learning

@@ Abstract
The recent surge in contrast-based graph self-supervised learning has prominently featured an intensified exploration of spectral cues. Spectral augmentation, which involves modifying a graph’s spectral properties such as eigenvalues or eigenvectors, is widely believed to enhance model performance. However, an intriguing paradox emerges, as methods grounded in seemingly conflicting assumptions regarding the spectral domain demonstrate notable enhancements in learning performance. Through extensive empirical studies, we find that simple edge perturbations - random edge dropping for node-level and random edge adding for graph-level self-supervised learning - consistently yield comparable or superior performance while being significantly more computationally efficient. This suggests that the computational overhead of sophisticated spectral augmentations may not justify their practical benefits. Our theoretical analysis of the InfoNCE loss bounds for shallow GNNs further supports this observation. The proposed insights represent a significant leap forward in the field, potentially refining the understanding and implementation of graph self-supervised learning.
@@ 1 Introduction
In recent years, graph learning has emerged as a powerhouse for handling complex data relationships in multiple fields, offering vast potential and value, particularly in domains such as data mining [12], computer vision [41], network analysis [6], and bioinformatics [15]. However, limited labels make graph learning challenging to apply in real-world scenarios. Inspired by the great success of Self-Supervised Learning (SSL) in other domains [8, 5], Graph Self-Supervised Learning (Graph SSL) has made rapid progress and has shown promise by achieving state-of-the-art performance on many tasks [40], where Contrast-based Graph SSL (CG-SSL) are most dominant [23]. This type of method is grounded in the concept of mutual information (MI) maximization. The primary goal is to maximize the estimated MI between augmented instances of the same object, such as nodes, subgraphs, or entire graphs. Among the new developments in CG-SSL, approaches inspired by graph spectral methods have garnered significant attention. A prevalent conviction is that spectral information, including the eigenvalues and eigenvectors of the graph’s Laplacian, plays a crucial role in enhancing the efficacy of CG-SSL [21, 17, 19, 42, 4].
@@
In general, methods in CG-SSL can be categorized into two types based on whether augmentation is performed on the input graph to generate different views [4]. i.e. augmentation-based and augmentation-free methods. Of the two, the augmentation-based methods are more predominant and widely studied [13, 23, 43, 21, 19, 42]. Specifically, spectral augmentation has received significant attention, as it modifies a graph’s spectral properties. This approach is believed to enhance model performance, aligning with the proposed importance of spectral information in CG-SSL. However, there seems no consensus on the true effectiveness of spectral information in the previous works proposing and studying spectral augmentation. SpCo [21] introduces the general graph augmentation (GAME) rule, which suggests that the difference in high-frequency parts between augmented graphs should be larger than that of low-frequency parts. SPAN [19] contends that effective topology augmentation should prioritize perturbing sensitive edges that have a substantial impact on the graph spectrum. Therefore, a principled augmentation method is designed by directly maximizing spectral change with a certain perturbation budget, without mentioning any specific domain of spectrum. GASSER [42] selectively perturbs graph structures based on spectral cues to better maintain the required invariance for contrastive learning frameworks. Specifically, it aims to augment the graphs to preserve task-relevant frequency components and perturb the task-irrelevant ones with care. While all three related methods are augmentation-based and share in the set of CG-SSL frameworks like GRACE [49] and MVGRL [13], a contradiction emerges among these related works on spectral augmentation: while SPAN advocates for maximizing the distance between the spectrum of augmented graphs regardless of spectral domains, SpCo and GASSER argue for the preservation of specific spectral components and domains during augmentation. The consistent performance gain derived from opposing methodical designs naturally raises our concern:
   •  Are spectral augmentations necessary in contrast-based graph SSL?
Given the question, this study aims to critically evaluate the effectiveness and significance of spectral augmentation in contrast-based graph SSL frameworks (CG-SSL). With evidence-supported claims and findings in the following sections, we show that despite their computational complexity, sophisticated spectral augmentations do not demonstrate clear advantages over simple edge perturbations. Our extensive experiments reveal that straightforward edge perturbations consistently achieve superior performance while being significantly more computationally efficient. Our theoretical analysis on the InfoNCE loss bounds for shallow GNNs provides additional insights into understanding this phenomenon and supports our claims. We elaborate on our findings through a series of studies carried out in the following efforts:
@@
    In Sec. 4, we demonstrate that shallow networks consistently achieve better performance in CG-SSL, analyze their inherent limitations in capturing global spectral information, and provide theoretical bounds on the InfoNCE loss that help explain the limited benefits of sophisticated spectral augmentations compared to simple edge perturbation.
@@
    In Sec 5, we claim that simple edge perturbation techniques, like adding edges to or dropping edges from the graph, not only compete well but often outperform spectral augmentations, without any significant help from spectral cues. To support this,
    (a) In Sec. 6, overall model performance on test accuracy with four state-of-the-art frameworks on both node- and graph-level classification tasks support the superiority of simple edge perturbation.
    (b) Studies in Sec. 7.1 reveal the indistinguishability between the average spectrum of augmented graphs from edge perturbation with optimal parameters on different datasets, no matter how different that of original graphs is, indicating GNN encoders can hardly learn spectral information from augmented graphs. That is to say, edge perturbations can not benefit from spectral information.
    (c) In Sec. 7.2, we analyze the effectiveness of state-of-the-art spectral augmentation baseline (i.e., SPAN) by perturbing edges to alter the spectral characteristics of augmented graphs from simple edge perturbation augmentation and examining the impact on model performance. As it turns out, the results show no performance degradation, indicating the spectral information contained in the augmentation is not significant to the model performance.
    (d) In Appendix E.3, statistical analysis is carried out to argue that the major reason edge perturbation works well is not because of the spectral information as they are statistically not the key factor on model performance.
@@ 2 Related work
Contrast-based Graph Self-Supervised (CG-SSL). CG-SSL learning alleviates the limitations of supervised learning, which heavily depends on labeled data and often suffers from limited generalization [22]. This makes it a promising approach for real-world applications where labeled data is scarce.
CG-SSL applies a variety of augmentations to the training graph to obtain augmented views. These augmented views, which are derived from the same original graph, are treated as positive sample pairs or sets. The key objective of CG-SSL is to maximize the mutual information between these views to learn robust and invariant representations. However, directly computing the mutual information of graph representations is challenging. Hence, in practice, CG-SSL frameworks aim to maximize the lower bound of mutual information using different estimators such as InfoNCE [11], Jensen-Shannon [26], and Donsker-Varadhan [1]. For instance, frameworks like GRACE [49], GCC [29], and GCA [50] utilize the InfoNCE estimator as their objective function. On the other hand, MVGRL [13] and InfoGraph [34] adopt the Jensen-Shannon estimator.
Some CG-SSL methods explore alternative principles. G-BT [2] extends the redundancy-reduction principle by decorrelating representations between two augmented views to prevent feature collapse. BGRL [35] adopts a momentum-driven Siamese architecture, using node feature masking and edge modification as augmentations to maximize mutual information between online and target network representations.
Graph Augmentations in CG-SSL. Beyond the choice of objective functions, another crucial aspect of augmentation-based methods in CG-SSL is the selection of augmentation techniques. Early work by [49] and [43] introduced several domain-agnostic heuristic graph augmentation for CG-SSL, such as edge perturbation, attribute masking, and subgraph sampling. These straightforward and effective methods have been widely adopted in subsequent CG-SSL frameworks due to their demonstrated success [35, 44]. However, these domain-agnostic graph augmentations often lack interpretability, making it difficult to understand the exact impact of these augmentations on the graph structure and learning outcomes.
To address this issue, MVGRL [13] introduces graph diffusion as an augmentation strategy, where the original graph provides local structural information and the diffused graph offers global context.
Moreover, three spectral augmentation methods–SpCo [21], GASSER [42], and SPAN [19]–stand out by offering design principles based on spectral graph theory, focusing on how to enhance CG-SSL performance through spectral manipulations.
However, our explorations show that these methods are unable to consistently outperform heuristic graph augmentations such as edge perturbation (DropEdge or AddEdge) in terms of performance under fair comparisons, and thus the design principles of graph augmentation still require further validation.
@@ 3 Preliminary study
Contrast-based graph self-supervised learning framework. CG-SSL captures invariant features of a graph by generating multiple views (typically two) through augmentations and then maximizing the mutual information between these views [40]. This approach is ultimately used to improve performance on various downstream tasks. Following previous work [39, 22, 40], we first denote the generic form of the augmentation $\mathcal{T}$ and objective functions $\mathcal{L}_{cl}$ of graph contrastive learning. Given a graph $\mathcal{G}=(\mathbf{A},\mathbf{X})$ with adjacency matrix $\mathbf{A}$ and feature matrix $\mathbf{X}$, the augmentation is defined as the transformation function $\mathcal{T}$. In this paper, we are mainly concerned with topological augmentation, in which feature matrix $\mathbf{X}$ remains intact:
@@
In this paper, we utilize four prominent CG-SSL frameworks to study the effect of spectral: MVGRL, GRACE, BGRL, and G-BT. MVGRL introduces graph diffusion as augmentation, while the other three frameworks use edge perturbation as augmentation. Each framework employs different strategies for its contrastive loss functions. MVGRL and GRACE use the Jensen-Shannon and InfoNCE estimators as object functions, respectively. In contrast, BGRL and G-BT adopt the BYOL loss [9] and Barlow Twins loss [45], which are designed to maximize the agreement between the augmented views without relying on negative samples. A more detailed description of the loss function can be found in the Appendix C.
Graph spectrum & Definition and application of spectral augmentation. We follow the standard definition of graph spectrum in this study, details of which can be found in Appendix B. Among various augmentation strategies proposed to enhance the robustness and generalization of graph neural networks, spectral augmentation has been considered a promising avenue [19, 21, 3, 42]. Spectral augmentation typically involves implicit modifications to the eigenvalues of the graph Laplacian, aiming at enhancing model performance by encouraging invariance to certain spectral properties. Among them, SPAN achieved state-of-the-art performance in both node classification and graph classification. In short, SPAN elaborates two augmentation functions, $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, where $\mathcal{T}_{1}$ maximizes the spectral norm in one view, and $\mathcal{T}_{2}$ minimizes it in the other view. Subsequently, these two augmentations are implemented in the four CG-SSL frameworks mentioned above (Strict definition in Appendix B). The paradigm used by SPAN aims to allow the GNN encoder to focus on robust spectral components and ignore the sensitive edges that can change the spectral drastically when perturbed.
@@ 4 Limitations of spectral augmentations
Limitations of shallow GNN encoders in capturing spectral information. Multiple previous studies indicate that shallow, rather than deep, GNN encoders can be effective in graph self-supervised learning. This might be the result of overfitting commonly witnessed in standard GNN tasks. We have also carried out many empirical studies with a range of CG-SSL frameworks and augmentations to support this idea in contrast-based graph SSL. As the most commonly applied GNN encoder in CG-SSL [43, 44, 10, 20], an empirical study on the relationship between the depth of GCN encoder and learning performance is conducted and results are presented in Fig. 1. From that, we can conclude that shallow GCN encoders with 1 or 2 layers usually have the best performance. Note that this tendency is not clear on graph-level tasks, which can be partially explained by the beneficial oversmoothing phenomenon present in this context [33]. It suggests that while deep encoders may have theoretically better expressive power than shallower ones, the limited benefits of deeper GNN architectures in the current CG-SSL practice imply that more layers may not bring significant improvements and could even hinder the quality of the learned graph representations.
@@
Figure 1: Accuracy of CG-SSL vs. number of GCN layers on node and graph classification on four datasets. (a) G-BT on node classification. (b) MVGRL on node classification. (c) G-BT on graph classification. (d) MVGRL on graph classification. We choose two representative datasets for each task, i.e. Cora and CiteSeer for the node-level and PROTEINS and IMDB-BINARY for the graph-level classification. The evaluation protocol, along with dataset details and other experimental settings, are provided in Section 6.1.
@@
By design, most GNN encoders primarily aggregate local neighborhood information through their layered structure, where each layer extends the receptive field by one hop. The depth of a GNN critically determines its ability to integrate information from various parts of the graph. With only a limited number of layers, a GNN’s receptive field is restricted to immediate neighborhoods (e.g., 1-hop or 2-hop distances). This limitation severely constrains the network’s ability to assimilate and leverage broader graph topologies or global features that are essential for encoding the spectral properties of the graph, given the definition of the graph spectrum.
@@
Limited implications for spectral augmentation in CG-SSL. Given the inherent limitations of shallow GNNs in capturing spectral information, the utility of spectral augmentation techniques in graph self-supervised learning settings warrants scrutiny. While spectral augmentation techniques modify the graph’s spectral components (e.g., eigenvalues and eigenvectors) to enrich the training process, their benefits may be limited if the primary encoder—a shallow GNN—cannot effectively process these spectral properties. To formally validate this intuition, we establish the following theoretical analysis.
Theoretical Analysis of InfoNCE Bounds. To better understand the effectiveness of augmentations in shallow GNNs, we derive the following theoretical bounds on InfoNCE loss:
Theorem 1 (InfoNCE Loss Bounds).
Given a graph $\mathcal{G}$ with minimum degree $d_{\min}$ and maximum degree $d_{\max}$, and its augmentation $\mathcal{G}^{\prime}$ with local topological perturbation strength $\delta$, for a $k$-layer GNN with ReLU activation and weight matrices satisfying $\left\|\mathbf{W}^{(l)}\right\|_{2}\leq L_{W}$, and assuming that the embeddings are normalized ($\left\|\mathbf{z}_{v}\right\|=\left\|\mathbf{z}^{\prime}_{v}\right\|=1$), the InfoNCE loss satisfies with high probability:
$-\log\left(\frac{e^{1/\tau}}{e^{1/\tau}+(n-1)e^{-\epsilon^{\prime}/\tau}}\right)\leq\mathcal{L}_{\text{InfoNCE}}(\mathcal{G},\mathcal{G}^{\prime})\leq-\log\left(\frac{e^{\left(1-\dfrac{\epsilon^{2}}{2}\right)/\tau}}{e^{\left(1-\dfrac{\epsilon^{2}}{2}\right)/\tau}+(n-1)e^{\epsilon^{\prime}/\tau}}\right),$ (3)
where $\epsilon$ is as defined in Lemma  5 and $\epsilon^{\prime}$ is as defined in Lemma  6. Detailed descriptions of notations can be found in Table  6. The proof of Theorem  1 and all related Lemmas can be found in Appendix  D.
This theoretical result reveals that the InfoNCE loss naturally stays within a narrow interval given a perturbation strength, regardless of augmentation complexity. Such a finding helps explain why sophisticated spectral augmentations may not significantly outperform simple ones in shallow architectures. While the potential benefits of coupling spectral augmentations with deeper GNN architectures remain an open question.
Numerical Estimation and Interpretation.
To illustrate the derived bounds, we provide a numerical estimation of the upper and lower bounds based on realistic parameters for 1-layer GNNs (which is the case for best performance for a bunch of benchmarks presented in Sec.  6 later). As detailed in Appendix  D.5, the parameters were chosen using typical graph augmentation settings and realistic assumptions about $\epsilon$ and $\epsilon^{\prime}$ derived from Lemma  5 and Lemma  6. The resulting bounds on the InfoNCE loss are: Lower bound: 4.7989, Upper bound: 5.4497. The difference between the two bounds is $5.4497-4.7989=0.6508$, indicating that the InfoNCE loss remains tightly constrained under these settings. This small interval suggests that shallow GNNs cannot fully utilize complex spectral augmentations, as their expressive capacity limits the potential variation in mutual information captured from augmented views. Our analysis reveals a critical insight: the limited efficacy of spectral augmentation stems from the inability of shallow GNNs to effectively capture and leverage the spectral properties of a graph. Instead, the learning outcomes are more directly influenced by simpler factors, such as the strength of edge perturbations. These findings reinforce the practicality of straightforward augmentation methods like edge dropping and adding, which perform comparably or better in this constrained theoretical setting.